OpenStudy (anonymous):

I'm trying to solve the equation: log base 3(x+2) + log base 3(4x-1)=5 and I have it to 0=4x^2+7x-245 and I can't figure out how to factor the trinomial. Can someone please help me?!?!

OpenStudy (anonymous):

oh 3^5 = 4x^2+7x-2

OpenStudy (anonymous):

because I ended up with: 3^5=4x^2+7x-2 so that went to 243=4x^2+7x-2 then I subtracted 243 from both sides and ended up with: 0=4x^2+7x-245

OpenStudy (anonymous):

thank you so much! I'll look at them and see if they help :D

OpenStudy (anonymous):

I listed the factors of 980 and followed the directions given in that second link to get the following: 28,35: 4x^2+7x-245 4x^2+35x-28x-245 x(4x+35)-7(4x-35) (4x+35)(x-7)

OpenStudy (anonymous):

ok that's what I was having trouble with :P I couldn't find the right factors

OpenStudy (anonymous):

x(4x+35)-7(4x+35)* oopsies

OpenStudy (anonymous):

tbh I had no idea how to factor things with leading coefficients until just now. You've learned me

OpenStudy (anonymous):

awesome both of us :D so could you help me finish out the equation to find what x is?

OpenStudy (anonymous):

\[\huge \log _3(x+2) + \log _3(4x-1)=5\]\[\huge \log _3(x+2)(4x-1)=5\] \[\huge \log _3(4x^2+7x-2)=5\]\[\huge (4x^2+7x-2)=3^5\]

OpenStudy (anonymous):

\[\huge (4x^2+7x-2)=243\]\[\huge 4x^2+7x-245=0\] \[\huge (4x+35)(x-7)=0\]

OpenStudy (anonymous):

 forgive my horrible handwriting